Circular No. 6089

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

PERIODIC COMET HARRINGTON (1994g) J. V. Scotti, Lunar and Planetary Laboratory, reports his discovery on

Oct. 4 Spacewatch images of two companion nuclei to this comet, separated from each other by 9′′ and from the principal nucleus (component A) by ∼ 23′ in p.a. 245^{o}. Given that the position of component A can be defined by T = 1994 Aug. 23.230 and the other elements on MPC 18258, components B and C correspond to T = 1994 Aug. 23.516 and 23.518, respectively. The total magnitudes of components A, B and C are 12.8, 21.3 and 20.2, respectively, and A and C show westward tails 10′ and 1′ long. Scotti confirmed the presence of the companions on Oct. 5 Spacewatch images obtained by R. Jedicke.

SUPERNOVAE 1994ab, 1994ac R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of two supernovae on U.K. Schmidt exposures by C. P. Cass and confirmed on 1.0-m reflector CCD images taken on Sept. 17 by G. J. Garradd and himself. SN 1994ab, discovered at mag 17 on a Sept. 13 I plate, is at α = 21^{h}13^{m}59^{s}.17, δ = −28^{o}32′26′′.3 (equinox 1950.0), 5′′.1 west and 16′′.8 south of the center of MCG −05-50-008; no image appears at the position on the ESO B or R surveys. SN 1994ac, discovered at mag 18.5 on a Sept. 11 R film, is at α = 23^{h}47^{m}19^{s}.91, δ = +2^{o}25′29′′.4 (equinox 1950.0), 3′′ west and 6′′.5 south of the center of UGC 12799 = MGC +00-60-052; no image appears at the position on U.K. Schmidt R and I plates from 1993 Sept. 12 and Oct. 21, respectively. Further confirmation of SN 1994ac is provided by an R film taken by M. Hartley on 1994 Oct. 3. A nearby star has end figures 17^{s}.42, 58′′.0.

PU VULPECULAE P. M. Garnavich, Harvard-Smithsonian Center for Astrophysics; and

S. R. Trammell, University of Chicago, communicate: “The current eclipse of the hot component in PU Vul has ended (cf. IAUC 5960, 5969). Spectra obtained on Oct. 3 using the McDonald Observatory’s 2.7-m telescope show that the broad emission lines of He II and C IV have returned. Monitoring at the Dominion Astrophysical Observatory constrains the onset of the eclipse to between 1993 Sept. 3 and 1994 Apr. 1. A spectrum taken on 1994 Aug. 1 showed no evidence of broad He II. The eclipse duration is thus between 5 and 13 months, and the orbital period of PU Vul is 13.6 ± 0.3 years (assuming that no other eclipses have occurred since 1980).”

1994 October 5 Brian G. Marsden